OpenStudy (anonymous):

Please help \[f(x)=\sqrt{x ^4-16x ^{2}}\] Let f be the function given by a) Find the domain of f. b) Find f '(x) c) Find the slope of the line normal to the graph of f at x = 5

OpenStudy (anonymous):

(x^4-16x^2)^(1/2) x^4 - 16x^2 >= 0 x^4 >= 16x^2 x^2 >= 16 x >= 4 so the domain is [4, infinity)

OpenStudy (anonymous):

derivative of (x^4-16x^2)^(1/2) (1/2)(x^4-16x^2)^(-1/2)(4x^3-32x) by the chain rule (4x^3-32x)/(2(sqrt(x^4-16x^2)))

OpenStudy (anonymous):

to find the slope at x=5, plug 5 into the equation above

OpenStudy (zarkon):

the domain is \[(-\infty,-4]\cup[0]\cup[4,\infty)\]

OpenStudy (anonymous):

Thanks and do you plug 5 into the original equation or f prime?

OpenStudy (anonymous):

plug it in to f prime to get the slope

OpenStudy (anonymous):

and then what? Sorry

OpenStudy (anonymous):

i got 34/3 for the slope

OpenStudy (anonymous):

that's it. the slope at x=5 is 34/3

OpenStudy (anonymous):

Oh alright thanks

OpenStudy (zarkon):

note that is just the slope of the tangent line...not the normal line

OpenStudy (anonymous):

How do you find the normal line then ?

OpenStudy (anonymous):

the reciprocal?

OpenStudy (zarkon):

if \(m\) is the slope of the tangent line then \(-\frac{1}{m}\) is the slope of the normal line.

OpenStudy (zarkon):

negative reciprocal

OpenStudy (anonymous):

so -3/34?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thanks